OpenStudy (anonymous):

-1 -5x+4 ----- = ---- x+2 x-5

OpenStudy (anonymous):

A. x=1,5 B. x=1 C. x=5 D. no solution

OpenStudy (nincompoop):

Solve for x over the real numbers: -1/(x+2) = (4-5 x)/(x-5) Write the left hand side as a single fraction. Bring -1/(x+2) together using the common denominator -x-2: 1/(-2-x) = (4-5 x)/(x-5) Multiply both sides by a polynomial to clear fractions. Cross multiply: x-5 = (4-5 x) (-2-x) Write the quadratic polynomial on the right hand side in standard form. Expand out terms of the right hand side: x-5 = 5 x^2+6 x-8 Move everything to the left hand side. Subtract 5 x^2+6 x-8 from both sides: -5 x^2-5 x+3 = 0 Write the quadratic equation in standard form. Divide both sides by -5: x^2+x-3/5 = 0 Solve the quadratic equation by completing the square. Add 3/5 to both sides: x^2+x = 3/5 Take one half of the coefficient of x and square it, then add it to both sides. Add 1/4 to both sides: x^2+x+1/4 = 17/20 Factor the left hand side. Write the left hand side as a square: (x+1/2)^2 = 17/20 Eliminate the exponent on the left hand side. Take the square root of both sides: x+1/2 = sqrt(17/5)/2 or x+1/2 = -sqrt(17/5)/2 Look at the first equation: Solve for x. Subtract 1/2 from both sides: x = sqrt(17/5)/2-1/2 or x+1/2 = -sqrt(17/5)/2 Look at the second equation: Solve for x. Subtract 1/2 from both sides: Answer: x = sqrt(17/5)/2-1/2 or x = -1/2-sqrt(17/5)/2

OpenStudy (anonymous):

so no solution?

OpenStudy (anonymous):

@nincompoop